A professor records the time (in minutes) that it takes 16 students to complete an exam. Compute the SS, the variance, and the standard deviation assuming the 16 students constitute a population and assuming the 16 students constitute a sample. (Round your answers for variance and standard deviation to two decimal places.)

23 44 42 29
41 20 25 43
40 52 39 36
50 25 14 33
Answer #1

From the given data :

Sample mean(xbar)=sumXi/n=556/16=34.75

Now the SS (sum of the squares)=sum(Xi-xbar)^2

Xbar(x-xbar)A2 23 34.75 138.0625 44 34.75 85.5625 42 34.75 52.5625 29 34.75 33.0625 41 34.75 39.0625 20 34.75 217.5625 25 34.75 95.0625 43 34.75 68.0625 40 34.75 27.5625 52 34.75 297.5625 39 34.75 18.0625 36 34.75 1.5625 50 34.75 232.5625 25 34.75 95.0625 14 34.75 430.5625 33 34.753.0625 (Sum)1835

So the SS (sum of the squares)=1835

Now the sample Variance = SS/(n-1)=1835/(16-1)=1835/15=122.33

Population variance=SS/n=1835/16=114.69

Sample standard deviation=√sample variance=√122.33=11.06

Population standard deviation=√population variance=10.71